A computer manufacturer can produce laptops at the rate of 100 per day. The manufacturer supplies its laptops to various computer retail outlets at a rate 65 per day. Set up cost for a production run is $500. Carrying cost is $125 per laptop a year. Assume the manufacturer operates 300 days a year. What is the optimal production run size?

  1. 39
  2. 395
  3. 667
  4. 828
  5. 612
Answer #1

1. C. 667

ANNUAL DEMAND = 19500(65 * 300)
SETUP COST = 500
HOLDING COST = 125
DAILY PRODUCTION = 100
DAILY USAGE = 65


EPQ = SQRT(2 * DEMAND * SETUP COST / HOLDING COST) * SQRT(DAILY PRODUCTION / DAILY PRODUCTION - DAILY USAGE)

EPQ = SQRT(2 * 19500 * 500 / 125) * SQRT(100 / 100 - 65) = 668